5. Write all the possible wave functions for two electrons in two generic orbitals 4, and 42
Answer #1

ANS:

For a 2- electron system, 2 characteristics should be considered:

1. electron's coordinates in wavefunctions must be such that the electrons are indistinguishable , i.e,

              |\small \psi2| =\small \psi . \small \psi as every electron is identical.

2.This requirement is according to pauli's exclusion principle. It is that the wavefunction for a 2 electron system must change sign any time the coordinates of any 2 electrons are permuted.

     \small \psi (1,2....i,..j...,N) = -\small \psi (1,2,...j,...i,....N)

Considering the above two requirements, the wavefunction for 2 electrons must be antisymmetric w.r.t interchange of spins and indistinguishable.

Thus, the possible wavefunctions are:

1\small \psi = 1/\small \sqrt{2(\small \phia(1)\small \phib(2) +\small \phib(1)\small \phia(2)) * 1/\small \sqrt{2}(\small \alpha 1\beta 2 - \beta 1\alpha 2)              Ms=0

3\small \psi=1/\small \sqrt{2}(\small \phia(1)\small \phib(2) - \small \phib(1)\small \phia(2)) *\tiny \begin{Bmatrix} \alpha 1\alpha 2 & & \\ 1/\sqrt{2}(\alpha 1\beta 2+\beta 1\alpha 2)& & \\ \beta 1\beta 2 & & \end{Bmatrix}              Ms =1, 0, -1